# Consciousness-Inspired Spatio-Temporal Abstraction for Better Generalization

### 04 Feb 2024 v3

@@ 3.4.1 Pruning
In this paper,we limit ourselves only to checkpoints from a return-unaware conditional generation model, leaving the question of how to improve the quality of the generated checkpoints for future work. Without learning, the proxy problem can be improved by making it more sparse, and making the proxy problem vertices more evenly spread in state space. To achieve this, we propose a pruning algorithm based on $k$-medoids clustering (Kaufman & Rousseeuw, 1990), which requires pairwise distance estimates between states. During proxy problem construction, we first sample a larger number of checkpoints, and then cluster them and select the centers (which are always real states).
@@
In addition to pruning the vertices, we also prune the edges according to a distance threshold, i.e., all edges with estimated distance over the threshold are deleted from the complete graph of the pruned vertices. This biases potential plans towards shorter-length, smaller-scale sub-problems, as far-away checkpoints are difficult for $\pi$ to achieve.
@@ 3.4.2 Safety & Delusion Control
Model-based HRL agents can be prone to blindly optimizing for objectives without understanding the consequences (Langosco et al., 2022; Paolo et al., 2022). We propose a technique to suppress delusions by exposing the edge estimators to potentially delusional targets that do not exist in the experience replay buffer. Details and examples are provided in the Appendix.
3.5 Overall Framework & Training
An intuitive overall algorithm and the training details are presented in Appendix H. The estimators are trained with KL-divergence and equal weighting of all terms, and the generator is trained with a VAE loss ,where the overall training loss is a simple sum of the aforementioned.
@@ 5 Experiments
As introduced in Sec. 2, our first goal is to test the zero-shot generalization ability of trained agents. To fully understand the results, it is necessary to have precise control of the difficulty of the training and evaluation tasks. Also, to validate if the empirical performance of our agents matches the formal analyses (Thm. 1), we need to know how close to the (optimal) ground truth our edge estimations and checkpoint policies are. These goals lead to the need for environments whose ground truth information (optimal policies, true distances between checkpoints, etc) can be computed. Thus, we base our experimental setting on the MiniGrid-BabyAI framework Chevalier-Boisvert et al. (2018b; a); Hui et al. (2020). Specifically, we build on the experiments used in our previous works Zhao et al. (2021); Alver & Precup (2022): the agent needs to navigate to the goal from its initial state in gridworlds filled with terminal lava traps generated randomly according to a difficulty parameter, which controls their density. During evaluation, the agent is always spawned at the opposite side from the goals. During training, the agent’s position is uniformly initialized to speed up training. We provide results for non-uniform training initialization in the Appendix.
These fully observable tasks prioritize on the challenge of reasoning over causal mechanisms over learning representations from complicated observations, which is not the focus of this work [^2] Skipper is implemented to be compatible with image inputs. The observations here are low-res two-channel (pseudo-)images ($12\times 12\times 2$). More details are presented in the Appendix.. Across all experiments, we sample training tasks from an environment distribution of difficulty $0.4$: each cell in the field has probability $0.4$ to be filled with lava while guaranteeing a path from the initial position to the goal. The evaluation tasks are sampled from increasing OOD difficulties of $0.25$, $0.35$, $0.45$ and $0.55$, where the training difficulty acts as a median. To step up the long(er) term generalization difficulty compared to existing work, we conduct experiments done on large, $12\times 12$ maze sizes, (see the visualization in Fig 2). The agents are trained for $1.5\times 10^{6}$ interactions.
@@
  • Skipper-once: A Skipper agent that generates one proxy problem at the start of the episode, and the replanning (choosing a checkpoint target based on the existing proxy problem) only triggers a quick selection of the immediate checkpoint target;
  • Skipper-regen: A Skipper agent that regenerates a proxy problem when replanning is triggered;
  • modelfree: A model-free baseline agent sharing the same base architecture with the Skipper variants, with a prioritized distributional Double DQN (Dabney et al., 2018; Van Hasselt et al., 2016);
@@ 5.1 Generalization Performance
Fig. 3 shows how the agents’ generalization performance evolves during training. These results are obtained with $50$ fixed sampled training tasks (different for each seed), a representative configuration of different numbers of training tasks including $\{1,5,25,50,100,\infty\}$333$\infty$ training tasks mean that an agent is trained on a different task for each episode. In reality, this may lead to prohibitive costs in creating the training environment., whose results are in the Appendix. In Fig. 3 a), we observe how well an agent performs on its training tasks. If an agent performs well here but badly in b), c), d) and e), e.g. the modelfree baseline, then we suspect that it overfitted on training tasks, likely an indicator of reliance on memorization (Cobbe et al., 2020).
We observe a (statistically-)significant advantage in the generalization performance of the Skipper agents throughout training. We have also included significance tests and power analyses (Colas et al., 2018; Patterson et al., 2023) in the Appendix, together with results for other training configurations. The regen variant exhibits dominating performance over all others. This is because the frequent reconstruction of the graph makes the agent less prone to being trapped in a low-quality proxy problem and provides extra adaptability in novel scenarios (more discussions in the Appendix). Skippers behave less optimally than expected during training, despite the strong generalization on evaluation tasks. As our ablation results and theoretical analyses consistently show, such a phenomenon is a composite outcome of inaccuracies both in the proxy problem and the checkpoint policy. One major symptom of an inaccurate proxy problem is that the agent would chase delusional targets. We address this behavior with a delusion suppression technique, to be discussed in the Appendix.
@@
Figure 3: Generalization Performance of Agents During Training: the $x$-axes correspond to training progress, while the aligned $y$-axes represent the success rate of episodes (optimal is 1.0). Each agent is trained with $50$ tasks. Each data point is the average success rate over $20$ evaluation episodes, and each error bar (95% confidence interval) is processed from $20$ independent seed runs. Training tasks performance is shown in (a) while OOD performance is shown in (b), (c), (d), (e).
Better than the modelfree baseline, LEAP obtains reasonable generalization performance, despite the extra budget it needs for pretraining. In the Appendix, we show that LEAP benefits largely from the delusion suppression technique. This indicates that optimizing for a path in the latent space is prone to errors caused by delusional subgoals. Lastly, we see that the Director agents suffer in these experiments despite their good performance in the single environment experimental settings reported by Hafner et al. (2022). We present additional experiments in the Appendix to show that Director is ill-suited for our generalization-focused setting: Director still performs well in single environment configurations, but its performance deteriorates fast with more training tasks. This indicates poor scalability in terms of generalization, a limitation to its application in real-world scenarios.
@@ 5.4 Summary of Experiments
Within the scope of the experiments, we conclude that the proposed framework provides benefits for generalization; And it achieves better generalization when exposed to more training tasks;
@@ 6 Conclusion & Future Work
Appendix A Experimental Results (Cont.)

### 16 Mar 2024 v4

@@ 3.4.1 Pruning
In this paper,we limit ourselves only to checkpoints from a return-unaware conditional generation model, leaving the question of how to improve the quality of the generated checkpoints for future work. Without learning, the proxy problem can be improved by making it more sparse, and making the proxy problem vertices more evenly spread in state space. To achieve this, we propose a pruning algorithm based on $k$-medoids clustering (Kaufman & Rousseeuw, 1990), which only requires pairwise distance estimates between states. During proxy problem construction, we first sample a larger number of checkpoints, and then cluster them and select the centers (which are always real states instead of imaginary weighted sums of state representations).
@@
In addition to pruning the vertices, we also prune the edges according to a distance threshold, i.e., all edges with estimated distance over the threshold are deleted from the complete graph of the pruned vertices. This biases potential plans towards shorter-length, smaller-scale sub-problems, as far-away checkpoints are difficult for $\pi$ to achieve, trading optimality for robustness.
@@ 3.4.2 Safety & Delusion Control
Model-based HRL agents can be prone to blindly optimizing for objectives without understanding the consequences (Langosco et al., 2022; Paolo et al., 2022). We propose a technique to suppress delusions by exposing edge estimation to potentially delusional targets that do not exist in the experience replay buffer. Details and examples are provided in the Appendix.
@@ 5 Experiments
As introduced in Sec. 2, our first goal is to test the zero-shot generalization ability of trained agents. To fully understand the results, it is necessary to have precise control of the difficulty of the training and evaluation tasks. Also, to validate if the empirical performance of our agents matches the formal analyses (Thm. 1), we need to know how close to the (optimal) ground truth our edge estimations and checkpoint policies are. These goals lead to the need for environments whose ground truth information (optimal policies, true distances between checkpoints, etc) can be computed. Thus, we base our experimental setting on the MiniGrid-BabyAI framework (Chevalier-Boisvert et al., 2018b; a; Hui et al., 2020). Specifically, we build on the experiments used in our previous works (Zhao et al., 2021; Alver & Precup, 2022): the agent needs to navigate to the goal from its initial state in gridworlds filled with terminal lava traps generated randomly according to a difficulty parameter, which controls their density. During evaluation, the agent is always spawned at the opposite side from the goals. During training, the agent’s position is uniformly initialized to speed up training. We provide results for non-uniform training initialization in the Appendix.
These fully observable tasks prioritize on the challenge of reasoning over causal mechanisms over learning representations from complicated observations, which is not the focus of this work. Across all experiments, we sample training tasks from an environment distribution of difficulty $0.4$: each cell in the field has probability $0.4$ to be filled with lava while guaranteeing a path from the initial position to the goal. The evaluation tasks are sampled from a gradient of OOD difficulties - $0.25$, $0.35$, $0.45$ and $0.55$, where the training difficulty acts as a median. To step up the long(er) term generalization difficulty compared to existing work, we conduct experiments done on large, $12\times 12$ maze sizes, (see the visualization in Fig 2). The agents are trained for $1.5\times 10^{6}$ interactions.
@@
  • Skipper-once: A Skipper agent that generates one proxy problem at the start of the episode, and the replanning (choosing a checkpoint target based on the existing proxy problem) only triggers a quick re-selection of the immediate checkpoint target;
  • Skipper-regen: A Skipper agent that re-generates a proxy problem when replanning is triggered;
  • modelfree: A model-free baseline agent sharing the same base architecture with the Skipper variants - a prioritized distributional Double DQN (Dabney et al., 2018; Van Hasselt et al., 2016);
@@ 5.1 Generalization Performance
Fig. 3 shows how the agents’ generalization performance evolves during training. These results are obtained with $50$ fixed sampled training tasks (different $50$s for each seed), a representative configuration of different numbers of training tasks including $\{1,5,25,50,100,\infty\}$222$\infty$ training tasks mean that an agent is trained on a different task for each episode. In reality, this may lead to prohibitive costs in creating the training environment., whose results are in the Appendix. In Fig. 3 a), we observe how well an agent performs on its training tasks. If an agent performs well here but badly in b), c), d) and e), e.g. the modelfree baseline, then we suspect that it overfitted on training tasks, likely indicating a reliance on memorization (Cobbe et al., 2020).
We observe a (statistically-)significant advantage in the generalization performance of the Skipper agents throughout training. We have also included significance tests and power analyses (Colas et al., 2018; Patterson et al., 2023) in the Appendix, together with results for other training configurations. The regen variant exhibits dominating performance over all others. This is likely due to the frequent reconstruction of the graph makes the agent less prone to being trapped in a low-quality proxy problem and provides extra adaptability in novel scenarios (more discussions in the Appendix). During training, Skippers behave less optimally than expected, despite the strong generalization on evaluation tasks. As our ablation results and theoretical analyses consistently show, such a phenomenon is a composite outcome of inaccuracies both in the proxy problem and the checkpoint policy. One major symptom of an inaccurate proxy problem is that the agent would chase delusional targets. We address this behavior with the delusion suppression technique, to be discussed in the Appendix.
@@
Figure 3: Generalization Performance of Agents During Training: the $x$-axes correspond to training progress, while the aligned $y$-axes represent the success rate of episodes (optimal is 1.0). Each agent is trained with $50$ tasks. Each data point is the average success rate over $20$ evaluation episodes, and each error bar (95% confidence interval) is processed from $20$ independent seed runs. Training tasks performance is shown in (a) while OOD evaluation performance is shown in (b), (c), (d), (e).
Better than the modelfree baseline, LEAP obtains reasonable generalization performance, despite the extra budget it needs for pretraining. In the Appendix, we show that LEAP benefits largely from the delusion suppression technique. This indicates that optimizing for a path in the latent space may be prone to errors caused by delusional subgoals. Lastly, we see that the Director agents suffer in these experiments despite their good performance in the single environment experimental settings reported by Hafner et al. (2022). We present additional experiments in the Appendix to show that Director is ill-suited for generalization-focused settings: Director still performs well in single environment configurations, but its performance deteriorates fast with more training tasks. This indicates poor scalability in terms of generalization, a limitation to its application in real-world scenarios.
@@ 5.4 Summary of Experiments
Within the scope of the experiments, we conclude that Skipper provides benefits for generalization; And it achieves better generalization when exposed to more training tasks;
@@ 6 Conclusion & Future Work
7 Reproducibility Statement
The results presented in the experiments are fully-reproducible with the open-sourced repository https://github.com/mila-iqia/Skipper.
8 Acknowledgements
Mingde is grateful for the financial support of the FRQ (Fonds de recherche du Québec) and the collaborative efforts during his time at Microsoft Research Montreal. We want to thank Riashat Islam, Ziyu Wang for the discussions in the initial stages of the project. We appreciate the computational resources allocated to us from Mila and McGill university.
Appendix A APPENDIX
Please use the following to quickly navigate to your points of interest.
  • Weaknesses & Limitations (Sec. B)
  • Skipper Algorithmic Details (Sec. C): pseudocodes, $k$-medoids based pruning, delusion suppression
  • Theoretical Analyses (Sec. D.1): detailed proofs, discussions
  • Implementation Details (Sec. E): for Skipper, LEAP and Director
  • More Experiments (Sec. F): experimental results that cannot be presented in the main paper due to page limit
  • Ablation Tests & Sensitivity Analyses (Sec. G)
Appendix B Weaknesses & Limitations
We would like to expand the discussions on the limitations to the current form of Skipper, as well as the design choices that we seek to improve in the future:
  • We generate future checkpoints at random by sampling the partial description space. Despite the post-processing such as pruning, the generated checkpoints do not prioritize on the predictable, important states that matter the most to form a meaningful long-term plan.
  • The current implementation is intended for pixel input fully-observable tasks with discrete state and action spaces. Such a minimalistic form is because we wish to isolate the unwanted challenges from other factors that are not closely related to the idea of this work, as well as to make the agent as generalist as possible. Skipper is naturally compatible with continuous actions spaces and the only thing we will need to do is to replace the baseline agent with a compatible one such as TD3 (Fujimoto et al., 2018); on the other hand, for continuous state spaces, the identification of the achievement of a checkpoint becomes tricky. This is due to the fact that a strict identity between the current state and the target checkpoint may be ever established, we either must adopt a distance measure for approximate state equivalence, or rely on the equivalence of the partial descriptions (which is adopted in the current implementation). We intentionally designed the partial descriptions to be in the form of bundles of binary variables, so that this comparison could be done fast and trivially for any forms of the state space; for partial observability, despite that no recurrent mechanism has been incorporated in the current implementation, the framework is not incompatible. To implement that, we will need to augment the state encoder with recurrent or memory mechanisms and we need to make the checkpoint generator directly work over the learned state representations. We acknowledge that future work is needed to verify Skipper’s performance on the popular partially-observable benchmark suites, which requires the incorporation of components to handle partial observability as well as scaling up the architectures for more expressive power;
  • We do not know the precise boundaries of the motivating theory on proxy problems, since it only indicates performance guarantees on the condition of estimation accuracy, which in turn does not correspond trivially to a set of well-defined problems. We are eager to explore, outside the scope of sparse-reward navigation, how this approach can be used to facilitate better generalization, and at the same time, try to find more powerful theories that guide us better;
Appendix C Skipper’s Algorithmic Details
C.1 Overall Skipper Framework (Pseudo-Code)
The pseudocode of Skipper is provided in Alg. 1, together with the hyperparameters used in our implementation.
for each episode do
       // — start of the subroutine to construct the proxy problem
       generate more than necessary (32) checkpoints by sampling from the partial descriptions given the extracted context from the initial state;
       $(k=12)$-medoid pruning upon estimated distances among all checkpoints; // prune vertices
       use estimators to annotate the edges between the nodes (including a terminal state estimator to correct the estimates);
       prune edges that are too far-fetched according to distance estimations (threshold set to be $8$, same as replan interval); // prune edges
       // — end of the subroutine to construct the proxy problem
       for each agent-environment interaction step until termination of episode do
             if decided to explore (DQN-style annealing $\epsilon$-greedy) then
                   take a random action;
            else
                         [OPTIONAL RE-GENERATION] call the subroutine above for Skipper-regen;
                         run value iteration (for $5$ iterations) on the proxy problem, select the target checkpoint;
                  follow the action suggested by the checkpoint-achieving policy;
            if time to train (every $4$ actions) then
                   sample hindsight transitions and train checkpoint-achieving policy, estimators (including a teriminal state estimator) and checkpoint generator;
                   [OPTIONAL DELUSION CONTROL]: train estimators using generated checkpoints;
            save interaction into the trajectory experience replay;
      convert trajectory into HER samples (relabel $4$ random states as additional goals);
C.2 k-medoids based pruning
We present the pseudocode of the modified $k$-medoids algorithm for pruning overcrowded checkpoints in Alg. 2. Note that the presented pseudocode is optimized for readers’ understanding, while the actual implementation is parallelized. The changes upon the original $k$-medoids algorithm is marked in purple, which implement a forced preservation of data points: when $k$-medoids is called after the unpruned graph is constructed, $\mathcal{S}_{\vee}$ is set to be the set containing the goal state only. This is intended to span more uniformly in the state space with checkpoints, while preserving the goal.
Let the estimated distance matrix be $D$, where each element $d_{i}j$ represents the estimated trajectory length it takes for $\pi$ to fulfill the transition from checkpoint $i$ to checkpoint $j$. Since $k$-medoids cannot handle infinite distances (e.g. from a terminal state to another state), the distance matrix $D$ is truncated, and then we take the elementwise minimum between the truncated $D$ and $D^{T}$ to preserve the one-way distances. The matrix containing the elementwise minimums would be the input of the pruning algorithm.
make sure $\mathcal{S}_{\vee}\subset\mathcal{S}_{\odot}$
repeat
       foreach medoid $M_{j}$ do
      Find the medoid $M_{j}$ with the lowest cost $J_{j}$;
       if $M_{j}$ changes then
             make sure $\mathcal{S}_{\vee}\subset\mathcal{S}_{\odot}$
             Replace $M_{j}$ with the data point in $C_{j}$ that minimizes the total cost;
until Convergence (no cost improvement);
Algorithm 2 Checkpoint Pruning with $k$-medoids
C.3 Delusion Suppression
RL agents are prone to blindly optimizing for an intrinsic objective without fully understanding the consequences of its actions. Particularly in model-based RL or in Hierarchical RL (HRL), there is a significant risk posed by the agents trying to achieve delusional future states that do not exist or beyond the safety constraints. With a use of a learned generative model, as in Skipper and other HP frameworks, such risk is almost inevitable, because of uncontrollable generalization effects.
Generalization abilities of the generative models are a double-edged sword. The agent would take advantage of its potentials to propose novel checkpoints to improve its behavior, but is also at risk of wanting to achieve non-existent unknown consequences. In Skipper, checkpoints imagined by the generative model could correspond to non-existent “states” that would lead to delusional edge estimates and therefore confuse planning. For instance, arbitrarily sampling partial descriptions may result in a delusional state where the agent is in a cell that can never be reached from the initial states. Since such states do not exist in the experience replay, the estimators will have not learned how to handle them appropriately when encountered in the generated proxy problem during decision time. We present a resulting failure mode in Fig. 5.
(a) generalization performance across training tasks
(b) distance estimation error (L1) for delusional edges
(c) frequency of selecting delusional targets
(d) target optimality
Figure 6: Performance of Skipper-once with the proposed Delusion Suppression Technique: each curve and corresponding error bar (95% CI) are processed from $20$ independent seed runs. a) the performance across training tasks is shown. A more optimal performance can be achieved with Skipper-once in training tasks, when delusions are suppressed; b) During training interactions, the error in estimated (truncated) distance from and to delusional targets are significantly reduced with the technique; c) The frequency of selecting a delusional target is reduced to almost negligible during the whole training process; d) The optimality of target checkpoint during training can be improved by the suppression. Each agent is trained with $50$ environments and each curve is processed from $20$ independent seed runs.
train the distance estimator $D$ as if these are sampled from the HER
Algorithm 3 Delusion Suppression
The delusion suppression technique can also be used to help us understand the failure modes of LEAP in Sec. E.2.4.
Appendix D Theoretical Analyses
D.1 Update Rules for Edge Estimation
First, we want to show that the update rules proposed in the main paper indeed estimate the desired cumulative discount and reward.
For the cumulative discounted reward random variable:
$=\gamma^{T_{\perp}-t}\mathbb{I}{\{S_{T_{\perp}}=s^{\odot}\}},$ (8)
D.2 Performance Bound
Similarly, we can derive a lower bound:
D.3 No Assumption on Optimality
Appendix E Implementation Details for Experiments
E.1 Skipper
E.1.1 Training
The agent is based on a distributional prioritized double DQN. All the trainable parameters are optimized with Adam at a rate of $2.5\times 10^{-4}$ (Kingma & Ba, 2014), with a gradient clipping by value (maximum absolute value $1.0$). The priorities for experience replay sampling are equal to the per-sample training loss.
E.1.2 Full State Encoder
E.1.3 Partial State Selector (Spatial Abstraction)
The selector $\sigma$ is implemented with one-head (not multiheaded, therefore the output linear transformation of the default multihead attention implementation in PyTorch is disabled.) top-$4$ attention, with each local perceptive field of size $8\times 8$ cells. Layer normalization (Ba et al., 2016) is used before and after the spatial abstraction.
E.1.4 Estimators
The estimators, which operate on the partial states, are $3$-layered MLPs with $256$ hidden units.
An additional estimator for termination is learned, which instead of taking a pair of partial states as input, takes only one, and is learned to classify terminal states with cross-entropy loss. The estimated distance from terminal states to other states would be overwritten with $\infty$. The internal $\gamma$ for intrinsic reward of $\pi$ is $0.95$, while the task $\gamma$ is $0.99$
The estimators use C51 distributional TD learning (Dabney et al., 2018). That is, the estimators output histograms (softmax over vector outputs) instead of scalars. We regress the histogram towards the targets, where these targets are skewed histograms of scalar values, towards which KL-divergence is used to train. At the output, there are $16$ bins for each histogram estimation (value for policy, reward, distance).
E.1.5 Recovering Discounts from Distances
We recover the distribution of the cumulative discount by replacing the support of the discretized truncated distances with the corresponding discounts, as illustrated in Fig. 7. This addresses the problem of $\mathbb{E}[\gamma^{D}]\neq\gamma^{\mathbb{E}[D]}$, as the probability of having a trajectory length of 4 under policy $\pi$ from state $s_{t}$ to $s_{\odot}$ is the same as a trajectory having discount $\gamma^{4}$.
E.1.6 Checkpoint Generator
Despite Skipper is designed to have the generator work on state level, that is, it should take learned state representations as inputs and have state representations as outputs, in our experiments, the generator actually operates on observation inputs and outputs. This is because of the preferred compactness of the observations and the equivalence to full states under full observability in our experiments.
The context extractor $\mathcal{E}_{c}$ is a $32$-dimensional BabyAI BOW embedder. It encodes an input observation into a representation of the episodic context.
The partial description extractor $\mathcal{E}_{z}$ is made of a $32$-dimensional BabyAI BOW embedder, followed by $3$ aforementioned residual blocks with $3\times 3$ convolutions (doubling the feature dimension every time) in between, ended by global maxpool and a final linear projection to the latent weights. The partial descriptions are bundles of $6$ binary latents, which could represent at most $64$ “kinds” of checkpoints. Inspired by VQ-VAE (Van Den Oord et al., 2017), we use the argmax of the latent weights as partial descriptions, instead of sampling according to the softmax-ed weights. This enables easy comparison of current state to the checkpoints in the partial description space, because each state deterministically corresponds to one partial description. We identify reaching a target checkpoint if the partial description of the current state matches that of the target.
The whole checkpoint generator is trained end-to-end with a standard VAE loss. That is the sum of a KL-divergence for the agent’s location, and the entropy of partial descriptions, weighted by $2.5\times 10^{-4}$, as suggested in https://github.com/AntixK/PyTorch-VAE. Note that the per-sample losses in the batches are not weighted for training according to priority from the experience replay.
We want to mention that if one does not want to generate non-goal terminal states as checkpoints, we could also seek to train on reversed $\langle S^{\odot},S_{t}\rangle$ pairs. In this case, the checkpoints to reconstruct will never be terminal.
E.1.7 HER
Each experienced transition is further duplicated into $4$ hindsight transitions at the end of each episode. Each of these transitions is combined with a randomly sampled observation from the same trajectory as the relabelled “goal”. The size of the hindsight buffer is extended to $4$ times that of the baseline that does not learn from hindsight accordingly, that is, $4\times 10^{6}$.
E.1.8 Planning
As introduced, we use value iteration over options (Sutton et al., 1999) to plan over the proxy problem represented as an SMDP. We use the matrix form $Q=R_{S\times S}+\Gamma V$, where $R$ and $\Gamma$ are the estimated edge matrices for cumulative rewards, respectively. Note that this notation is different from the ones we used in the manuscript. The checkpoint value $V$, initialized as all-zero, is taken on the maximum of $Q$ along the checkpoint target (the actions for $\mu$) dimension. When planning is initiated during decision time, the value iteration step is called $5$ times. We do not run until convergence since with low-quality estimates during the early stages of the learning, this would be a waste of time. The edges from the current state towards other states are always set to be one-directional, and the self-loops are also removed. This means the first column as well as the diagonal elements of $R$ and $\Gamma$ are all zeros. Besides pruning edges based on the distance threshold, as introduced in the main paper, the terminal estimator is also used to prune the matrices $R$ and $\Gamma$: the rows corresponding to the terminal states are all zeros.
E.1.9 Hyperparameter Tuning
Some hyperparameters introduced by Skipper can be located in the pseudocode in Alg. 1.
This leads to how we tuned the pruning (distance) threshold: we fully used the advantage of our experiments on DP-solvable tasks: with a snapshot of the agent during its training, we can sample many $\langle$ starting state, target state $\rangle$ pairs and calculate the ground truth distance between the pair, as well as the failure rate of reaching from the starting state to the target state given the current policy $\pi$, then plot them as the $x$ and $y$ values respectively for visualization. We found such curves to evolve from high failure rate at the beginning, to a monotonically increasing curve, where at small true distances, the failure rates are near zero. We picked $8$ because the curve starts to grow explosively when the true distances are more than $9$.
$k$ for $k$-medoids We tuned this by running a sensitivity analysis on Skipper agents with different $k$’s, whose results are presented previously in this Appendix.
E.2 LEAP
E.2.1 Adaptation for Discrete Action Spaces
The LEAP baseline has been implemented from scratch for our experiments, since the original open-sourced implementation [^4] https://github.com/snasiriany/leap was not compatible with environments with discrete action spaces. LEAP’s training involves two pretraining stages, that are, generator pretraining and distance estimator pretraining, which were originally named the VAE and RL pretrainings. Despite our best effort, that is to be covered in detail, we found that LEAP was unable to get a reasonable performance in its original form after rebasing it on a discrete model-free RL baseline.
E.2.2 Replacing the Model
We tried to identify the reasons why the generalization performance of the adapted LEAP was unsatisfactory: we found that the original VAE used in LEAP is not capable to handle even few training tasks, let alone generalize well to the evaluation tasks. Even by combining the idea of the context / partial description split (still with continuous latents), during decision time, the planning results given by the evolutionary algorithm (Cross Entropy Method, CEM, Rubinstein (1997a)) almost always produce delusional plans that are catastrophic in terms of performance. This was why we switched into LEAP the same conditional generator we proposed in the paper, and adapted CEM accordingly, due to the change from continuous latents to discrete.
We also did not find that using the pretrained VAE representation as the state representation during the second stage helped the agent’s performance, as the paper claimed. In fact, the adapted LEAP variant could only achieve decent performance after learning a state representation from scratch in the RL pretraining phase. Adopting Skipper’s splitting generator also disables such choice.
E.2.3 Replacing TDM
The original distance estimator based on Temporal Difference Models (TDM) also does not show capable performance in estimating the length of trajectories, even with the help of a ground truth distance function (calculated with DP). Therefore, we switched to learning the distance estimates with our proposed method. Our distance estimator is not sensitive to the sub-goal time budget as TDM and is hence more versatile in environments like that was used in the main paper, where the trajectory length of each checkpoint transition could highly vary. Like for Skipper, an additional terminal estimator has been learned to make LEAP planning compatible with the terminal lava states. Note that this LEAP variant was trained on the same sampling scheme with HER as in Skipper.
E.2.4 Failure Mode: Delusional Plans
(a) training, diff $0.4$
(b) OOD, diff $0.25$
(c) OOD, diff $0.35$
(d) OOD, diff $0.45$
(e) OOD, diff $0.55$
E.3 Director
Figure 9: An example for simplified observations for Director.
E.3.1 Adaptation
We based our experiments of Director (Hafner et al., 2022) on the publicly available code (https://github.com/danijar/director) released by the authors. Except for a few changes in the parameters, which are depicted in Tab. 1, we have used the default configuration provided for Atari environments. Note that as the Director version in which the worker receives no task rewards performed worse in our tasks, we have used the version in which the worker receives scaled task rewards (referred to as “Director (worker task reward)” in Hafner et al. (2022)). This agent has also been shown to perform better across various domains in Hafner et al. (2022).
Encoder. Unlike Skipper and LEAP agents, the Director agent receives as input a simplified RGB image of the current state of the environment (see Fig. 9). This is because we found that Director performed better with its original architecture, which was designed for image-based observations. We removed all textures to simplify the RGB observations.
Parameter
Value
replay_size
2M
replay_chunk
12
imag_horizon
8
env_skill_duration
4
train_skill_duration
4
worker_rews
{extr: 0.5, expl: 0.0, goal: 1.0}
sticky
False
gray
False
E.3.2 Failure Modes: Bad Generalization, Sensitive to Short Trajectories
Training Performance. We investigated why Director is unable to achieve good training performance(Fig. 3). As Director was designed to be trained solely on environments where it is able to collect long trajectories to train a good enough recurrent world model (Hafner et al., 2022), we hypothesized that Director may perform better in domains where it is able to interact with the environment through longer trajectories by having better recurrent world models (i.e., the agent does not immediately die as a result of interacting with specific objects in the environment). To test this, we experimented with variants of the used tasks, where the lava cells are replaced with wall cells, so the agent does not die upon trying to move towards them (we refer to this environment as the “walled” environment). The corresponding results on $50$ training tasks are depicted in Fig. 10. As can be seen, the Director agent indeed performs better within the training tasks than in the environments with lava.
Generalization Performance. We also investigated why Director is unable to achieve good generalization (Fig. 3). As Director trains its policies solely from the imagined trajectories predicted by its learned world model, we believe that the low generalization performance is due to Director being unable to learn a good enough world model that generalizes to the evaluation tasks. The generalization performances in both the “walled” and regular environments, depicted in Fig. 10, indeed support this argument. Similar to what we did in the main paper, we also present experimental results for how the generalization performance changes with the number of training environments. Results in Fig. 11 show that the number of training environments has little effect on its poor generalization performance.
(a) training, diff $0.4$
(b) OOD, diff $0.25$
(c) OOD, diff $0.35$
(d) OOD, diff $0.45$
(e) OOD, diff $0.55$
(a) training, diff $0.4$
(b) OOD, diff $0.25$
(c) OOD, diff $0.35$
(d) OOD, diff $0.45$
(e) OOD, diff $0.55$
Appendix F Experimental Results (Cont.)